OpenStudy (anonymous):

Differentiate y = 3x^2 – 5sin3x

myininaya (myininaya):

\[y'=(3x^2-5 \sin(3x))'=(3x^2)'-(5 \sin(3x))'=3(x^2)'-5(\sin(3x))'\] \[=3(2x)-5(3 \cos(3x))=6x-15 \cos(3x)\]

OpenStudy (anonymous):

THANK YOU!!!!!

myininaya (myininaya):

np i hope you enjoy the many steps i gave you lol

OpenStudy (anonymous):

I do. I need them too